Operations Report – January 14th

SOL: 2

Name of person filing report: Freddy Castaneda

Non-nominal systems: Suit #9 still not working, Suits #3, 6 and 8 under charging review, a Suit charger is needed because there is one that does not work and no spares are available.

Notes on non-nominal systems: Suit #9 was repaired, the electrical connections were fixed and a broken 5A fuse was identified but it could not be changed because there is no spare fuses, everything but the fuse is operational in the suit at this moment. Suits #3, 6 and 8 are under charging review, pending for diagnostic.

A Suit charger and spare ones are needed.

Generator (hours run): 17hr 37min; Turned on last night (13Jan2019) at 17:40; Turned off this morning (14Jan2019) at 11:17 cloudy day…; Turned on tonight (14Jan2019) at 16:45

Solar SOC – Turned on (13Jan2019) 85%; Turned off (14Jan2019) 100%; Turned on (14Jan2019) 88%

Diesel Reading – Just under 50%

Propane Reading – Under 30%, almost 29%

Ethanol Free Gasoline – Not in use

Water (auxiliary tank) – Not in use

Water (static tank) – About 90%; 480 gallons

Auxiliary to Static tank transfer – No

Gallons transferred: Not applicable

Water in GreenHab – About 87%; 260 gallons

Water (loft) – At level marker 10

Static to Loft Pump used – Yes; At 8:45 to refill tank

Water Meter: 01400658

Toilet tank emptied: Yes, a few minutes ago and still feels the smell 😉

Deimos rover used: No, still not functional

Hours: Not applicable

Beginning charge: Not applicable

Ending charge: Not applicable

Currently charging: Not applicable

Sojourner rover used: ASSIGNED TO DIRECTOR

Hours: Not applicable

Beginning charge: Not applicable

Ending charge: Not applicable

Currently charging: Not applicable

Spirit rover used: Yes, used for training EVA´s 1 and 2

Hours: 67.7

Beginning charge: 100%

Ending charge: 90%

Currently charging: Yes

Opportunity rover used: Not used

Hours: 45.9

Beginning charge: 100%

Ending charge: 100%

Currently charging: Yes

Curiosity rover used: Yes, used for training EVA’s 1 and 2

Hours: 69.7

Beginning charge: 100%

Ending charge: 95%

Currently charging: Yes

Notes on rovers: Brakes on Opportunity still not working properly

ATV’s Used: None (Honda, 300, 350.1, 350.2, 350.3)

Reason for use: None

Oil Added? No

ATV Fuel Used: None

# Hours the ATVs were Used today: None

Notes on ATVs: Nothing to report

HabCar used and why, where? Not used

CrewCar used and why, where? Refueled and returned by the Crew 202 with 172540 miles

General notes and comments: Nothing to report

Summary of internet: Nothing to report

Summary of suits and radios: Suit #9 has a broken 5A fuse, still not working, Suits 3,6 and 8 under charging review, one charger is not working; Suit #7 has not been reviewed yet; I´m aware to make a diagnosis.

Summary of Hab operations: Some crew members after a hard review, based on their skills on different engineering areas, have identified that the gas-like smell is probably generated by the metal of the stairs, the welding residues and the oxide, we recommend to give a new layer of paint to the staircase to decrease the smell and protect the metal which is showing some signs of corrosion.

Summary of GreenHab operations: Nothing to report.

Summary of ScienceDome operations: Nothing to report.

Summary of RAM operations: We were informed this morning by Shannon that yesterday the power was tripped to the power stations because the rover plugs were not completely clean before they were plugged back into power, so the rovers were not charged, she fixed it this morning.

Summary of any observatory issues: Nothing to report.

Summary of health and safety issues: Low batteries on two smoke detectors, spare batteries are needed.

Questions, concerns and requests to Mission Support: Nothing to report